An airline has 6 flights that it wishes to advertise in one-page ads in a sunday newspaper. It wants to feature a different flight each sunday for six sundays. A) how many different sequences of ads can the airline run? B) suppose that the airline decides at the last minute to run only 4 ads. how many sequences of ads can it run? show working please.
Answer by robertb(5830)

A) 6! = 6*5*4*3*2*1= 720 sequences.
B) 6P4 = 6*5*4*3 = 360 sequences.